2x +3y=10 , y=5
j,ildyfailysgdbvisdygvw9duvbu9oebvbsdljkmnhbj bvhyyhbcdffhgbvfcgyhbvgbhyhuhbvfcgtyuhuhvfcgvgfrtynbvfgrt

Grade:10

1 Answers

Nikhil
3 years ago
2x + 3(5) = 10
2x + 15 = 10
2x = 10 -15
2x = -5
x = -5/2
The value of x in given equation is -5/2 or -2.5.